OpenStudy (vortish):

the perimeter of a rectangle is 36 ft the length is 2 ft longer than the width / find the solution write a system of linear equations and solve the resulting system 2x+2y = 2l +2w correct

OpenStudy (anonymous):

l = length w = width Perimeter = 36 perimeter = 2l + 2w l = w + 2 substitute 2(w+2) + 2w = 36 2w + 4 + 2w = 36 4w + 4 = 36 4w = 32 w = 8 l = 8 + 2 l = 10